OpenStudy (anonymous):

Find the exact value of the composition. @ganeshie8

OpenStudy (anonymous):

is there a way to put it in my calculater?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

can anyone help me please!

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

first find \(\large \sin (\frac{\pi}{2})\)